Identify the ordered pair(s) so that it (they) is (are) not a solution(s) for 7x + 2y = 17:
(0, 2/17)
(1, 17/7)
(0.5, 27/4)
(-2, 31/2)

Answer by checkley71(8403)

7X+2Y=17 SUBSTITUTING THE X & Y VALUES WE'LL SEE IF THEY=17.
7*0+2*2/17=17
0+4/17=17
4/17=17 NOT A SOLUTION
7*1+2*17/7=17
7+24/7=17
7+3.428=17
10.328=17 NOT AN ANSWER.
7*.5+2*27/4=17
3.5+54/4=17
3.5+13.5=17
17=17 THIS LOOKS LIKE A SOLUTION (.5, 27/4)
7*-2+2*31/2=17
-14+62/2=17
-14+31=17
17=17 THIS ALSO LOOKS GOOD. (-2,31/2)